s. If you increase the temperature of a volume of air by four times while at the same time doubling the pressure, would the volume of the air increase or decrease? Why? continue 1464B
Answer #1

Suppose the initial pressure , volume and temperature is P1, V1 and T1

Final pressure, volume and temperature is P2, V2 and T2.

P1V1/T1 = P2V2/T2

T2 = 4T1 and P2 = 2P1

P1V1/T1 = 2P1 * V2/4T1

or, V1 = V2/2

So, the volume will also increase. It will be double of the initial volume.